OpenStudy (anonymous):

-x-4>3x-2 (Solve the inequality)

OpenStudy (shadowfiend):

First things first, collect like terms on either side: \begin{align} -x - 4 &> 3x - 2\\ -x &> 3x - 2 + 4\\ -x - 3x &> -2 + 4\\ -4x &> 2 \end{align}

OpenStudy (shadowfiend):

Once we've done that, we can divide to get the x alone, but! When you divide by a negative and you have an inequality, you have to flip the direction of the inequality! So: \begin{align} -4x &> 2\\ x &< -\frac{1}{2} \end{align}

OpenStudy (anonymous):

so this somewat has to do with standeredform/liketerms

OpenStudy (shadowfiend):

Now, let's check it and make sure I didn't mess up. We'll try x = -1/2, which should *not* satisfy the inequality. x = -1/2 means: \begin{align} -\frac{1}{2} - 4 &> 3\frac{-1}{2} - 2\\ -\frac{9}{2} &> -\frac{3}{2} - 2\\ -\frac{9}{2} &> -\frac{7}{2} \end{align} This isn't quite true, which is what we expect. Now let's try something less than -1/2, namely -1: \begin{align} -1 - 4 &> 3(-1) - 2\\ -5 &> -4 - 2\\ -5 &> -6 \end{align} That's right! So it looks like the solution makes sense.

OpenStudy (shadowfiend):

Yes. Solving inequalities is literally the same thing as solving a regular equation, except that when you multiply or divide by a negative, you have to flip the inequality sign.